OpenStudy (christos):

Hexadecimal, Express the hexadecimal 1A2F16 in decimal.

OpenStudy (anonymous):

decimal = 1715990

OpenStudy (christos):

How did you do it?

OpenStudy (anonymous):

i googled hexadecimal to decimal and clicked on the first link: http://easycalculation.com/hex-converter.php it has a nifty little table if you want to do it yourself

OpenStudy (anonymous):

DEC 0 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 HEX 0 1 2 3 4 5 6 7 8 9 A B C D E F

OpenStudy (mathstudent55):

Hexadecimal numbers work like decimal numbers with the exception that hexadecimal digits go from 0 to 15 instead of 0 to 9. Decimal number: 987,654,321 1 - ones 2 - tens 3 - hundreds 4 - thousands etc. each place to the left is 10 times larger than the previous one. hexadecimal number: 987,654,321 1 - how many "units" from 0 to 15 2 - how many 16's 3 - how many 256's or 16^2's 4 - how many 4096's or 16^3's etc. each place to the left is 16 times larger than the previous one. Use the table that Chanellos wrote above for each digit. Starting from the right-most digit 1A2F16 = 6 * 1 + 1 * 16 + 15 * 16^2 + 2 * 16^3 + 10 * 16^4 + 1 * 16^5 = 1,715,990

OpenStudy (christos):

So according to you guys this one is a mistake? http://screencast.com/t/WdhgaZlu6

OpenStudy (anonymous):

what does that one get? it won't open for me

hartnn (hartnn):

that is actually \(\huge(1A2F)_{16}\) we thought its (1A2F16) that is correct

OpenStudy (christos):

How did you come up to this conclusion hartnn?

hartnn (hartnn):

because in your image 16 is not considered as a part of the number. it is just used to denote that the base of the number is 16 but the way its written, we initially thought that the number is 1A2F16

OpenStudy (anonymous):

subscripts on numbers are sometimes used to denote that the number is in a base other than 10.

OpenStudy (christos):

http://screencast.com/t/8WYOy9RetX2x but here it is used like number as the rest of the digits?

OpenStudy (anonymous):

I can't see that.

OpenStudy (anonymous):

maybe it's that I'm in China

OpenStudy (christos):

nvm firget abiut it I got it but, the solution I come up to if I convert "1A2F" to decimal is 6703 ?

hartnn (hartnn):

isn't it the same same image ? and no, the 16 is not used like number. the digit at unit's place is F which = 15 in decimal. and that was multiplied by 16^0=1 hmmm...ok

hartnn (hartnn):

yes, in decimal its 6703

OpenStudy (christos):

http://screencast.com/t/7qzQhBfATA

OpenStudy (anonymous):

yes, 6703

hartnn (hartnn):

if you want to specify base, you can write, \(\huge (1A2F)_{16}=(6703)_{10}\)

OpenStudy (christos):

oooh

hartnn (hartnn):

either write 6703 only or \((6703)_{10}\) 670310 is incorrect. i prefer 2nd.

OpenStudy (mathstudent55):

Now that we actually understand what the question is thanks to hartnn, here we go again: Use the table that Chanellos wrote above for each digit. Starting from the right-most digit (1A2F)base16 = = F * 1 + 2 * 16 + A * 16^2 + 1 * 16^3 = 15 * 1 + 2 * 16 + 10 * 16^2 + 1 * 16^3 = (6703)base10